# Smooth Tests for Normality in ANOVA

### v3

@@ 4 Data-driven choice of $K$
The methodology in Section 3 relies on the fact that the order of smooth alternatives (2.1), $K$, is fixed before testing. The choice of $K$ critically affects the performance of smooth tests (Ledwina,, 1994; Kallenberg and Ledwina, 1995a, ; Kallenberg and Ledwina, 1995b, ; Inglot and Ledwina,, 1996; Inglot et al.,, 1997; Kallenberg and Ledwina,, 1997; Kallenberg and Teresa,, 1997; Kallenberg and Ledwina,, 1999; Janic-Wróblewska and Ledwina,, 2000; Ducharme and Fontez,, 2004; Ducharme and Lafaye de Micheaux,, 2004; Kraus,, 2007; Duchesne et al.,, 2016). A large $K$ introduces redundant components in the alternatives that contribute little to the test statistic but inflate the degrees of freedom, leading to power dilution. Conversely, a small $K$ increases the risk that $\mathrm{E}[\pi_{k}(Z)]=0$ for all $1\leq k\leq K$, rendering the test powerless. Hence, an appropriate choice of $K$ is crucial. In this section, we focus on data-driven selection of $K$ for the three cases in Section 3.
In the literature of Neyman’s smooth tests, data-driven selection was first proposed by Ledwina, (1994) for testing uniformity. The data-driven procedure consists of two steps. First, Schwarz’s selection rule (also known as the Bayesian Information Criterion, BIC) is applied to determine a suitable dimension for the smooth model that best fits the data. Second, Neyman’s smooth test is performed within the selected model space, yielding a data-driven test statistic. The data-driven smooth test thus serves to combine preliminary model selection with a more precise inferential procedure, being Neyman’s test in the “right” direction. Desirable theoretical properties and extensive numerical results of the selection rule and the induced test statistic were established in Kallenberg and Ledwina, 1995a and Kallenberg and Ledwina, 1995b , Inglot and Ledwina, (1996). Later, modifications of Schwarz’s rule addressed more complex problems, including testing for composite hypotheses (Kallenberg and Ledwina,, 1997; Kallenberg and Teresa,, 1997; Inglot et al.,, 1997), testing for independence (Kallenberg and Ledwina,, 1999), two-sample testing problems (Janic-Wróblewska and Ledwina,, 2000), and so on. These modified rules leverage smooth test statistics directly, avoiding the computation of the maximized log-likelihood, and are easier to implement.
@@
$\hat{K}=\min\left\{\mathop{\arg\max}\limits_{1\leq k\leq D}\left(N\hat{\Psi}_{k}^{2}-k\log N\right)\right\},$ (4.1)
where $D$ is a fixed positive integer, and $\hat{\Psi}_{k}^{2}$ takes the form of (3.3) or (3.12) depending on the specific scenario. The resulting data-driven test statistic is $N\Psi_{\hat{K}}^{2}$, with $\hat{K}$ given by (4.1). The form of (4.1) is mainly inspired by Ducharme and Fontez, (2004), Ducharme and Lafaye de Micheaux, (2004), Kraus, (2007) and Duchesne et al., (2016) where the upper bound of selection $D$ is fixed. An alternative is to let $D=D(N)\to\infty$ as $N\to\infty$ (see, for example, Kallenberg and Ledwina, 1995a, ; Kallenberg and Ledwina, 1995b, ; Inglot and Ledwina,, 1996; Inglot et al.,, 1997; Kallenberg and Ledwina,, 1997; Kallenberg and Teresa,, 1997; Kallenberg and Ledwina,, 1999; Janic-Wróblewska and Ledwina,, 2000). Although a diverging upper bound can theoretically improve the consistency of smooth tests, the required rate of divergence is slow, and simulation studies show that empirical power levels off rapidly as $D$ increases. Hence, our proposed selection rule (4.1) with fixed $D$ is both reasonable and easy to implement in practice.
To establish the theoretical properties of $\hat{K}$ and $N\hat{\Psi}_{\hat{K}}^{2}$, we introduce a revised version of the alternative hypothesis $H_{1}$ (3.4) and Assumption 3.2. Consider
$H_{1}^{\prime}:\mathrm{E}[\pi_{1}(Z)]=\ldots=\mathrm{E}[\pi_{K_{0}-1}(Z)]=0,\mathrm{E}[\pi_{K_{0}}(Z)]\neq 0,\quad K_{0}\leq D,$ (4.2)
@@
    For model (3.1) in Section 3.1, as $N\to\infty$, under $H_{0}:e\sim\Phi(x)$, $\mathrm{P}(\hat{K}=1)\to 1$ and $N\hat{\Psi}_{\hat{K}}^{2}\to_{d}\chi_{1}^{2}$; under $H_{1}^{\prime}$ in (4.2), $\mathrm{P}(\hat{K}\geq K_{0})\to 1$ and $\mathrm{P}(N\hat{\Psi}^{2}_{\hat{K}}\leq x)\to 0$ for any $x\in\mathbb{R}$.
@@
    For model (3.9) in Section 3.2, as $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$, and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$, under $H_{0}:e\sim\Phi(x)$, $\mathrm{P}(\hat{K}=1)\to 1$ and $N\hat{\Psi}_{\hat{K}}^{2}\to_{d}\chi_{1}^{2}$; under $H_{1}^{\prime}$ in (4.2), $\mathrm{P}(\hat{K}\geq K_{0})\to 1$ and $\mathrm{P}(N\hat{\Psi}^{2}_{\hat{K}}\leq x)\to 0$ for any $x\in\mathbb{R}$.
@@
    For model (3.10) in Section 3.3, with additional conditions such that (i) there exist $0<\underline{\sigma}\leq\overline{\sigma}<\infty$ such that $\underline{\sigma}<\inf_{1\leq j\leq J}\sigma_{j}\leq\sup_{1\leq j\leq J}\sigma_{j}<\overline{\sigma}$; (ii) there exist $0<\underline{q}\leq\overline{q}<\infty$ such that $\underline{q}<\inf_{1\leq j\leq J}q_{j}\leq\sup_{1\leq j\leq J}q_{j}<\overline{q}$, as $\min\{N_{1},\ldots,N_{J}\}\to\infty$ and $J=o(N^{1/2})$, under $H_{0}:e\sim\Phi(x)$, $\mathrm{P}(\hat{K}=1)\to 1$, $N\hat{\Psi}_{\hat{K}}^{2}\to_{d}\chi_{1}^{2}$ and $N\tilde{\Psi}_{\hat{K}}^{2}\to_{d}\chi_{1}^{2}$; under $H_{1}^{\prime}$ in (4.2), $\mathrm{P}(\hat{K}\geq K_{0})\to 1$ and $\mathrm{P}(N\tilde{\Psi}^{2}_{\hat{K}}\leq x)\to 0$ for any $x\in\mathbb{R}$.
Theorem 10 presents the unified results of data-driven smooth tests for the three cases in Section 3. Under the null hypothesis, the probability of $\{\hat{K}=1\}$ tends to $1$ asymptotically, implying that the first-order smooth model provides the best fit to $\{\hat{Z}_{ij}\}_{i=1,j=1}^{N_{j},J}$ among the $D$ candidate models. From the perspective of hypothesis testing, this means that $\hat{\Psi}_{1}^{2}$ is informative and sufficient for assessing uniformity or normality. Meanwhile, the data-driven selection procedure allows for the identification of the mechanism behind nonuniformity (and also nonnormality) and enhances the test’s power against a broader class of alternatives, $H_{1}^{\prime}$. These findings are consistent with previous results on data-driven smooth tests.
In practice, however, the $\chi_{1}^{2}$ limiting null distribution of data-driven smooth test statistics often performs poorly in finite samples (Kallenberg and Ledwina, 1995b, ; Inglot et al.,, 1997; Kallenberg and Ledwina,, 1997; Kallenberg and Teresa,, 1997; Kallenberg and Ledwina,, 1999; Janic-Wróblewska and Ledwina,, 2000; Ducharme and Fontez,, 2004; Ducharme and Lafaye de Micheaux,, 2004; Kraus,, 2007; Duchesne et al.,, 2016). To address this issue in our problem, and provided the nested orthonormal system (i.e. $\{\pi_{k}\}_{k=1}^{K}\subseteq\{\pi_{k}\}_{k=1}^{K+1}$ for $1\leq K\leq D-1$), we follow the approach of Kallenberg and Ledwina, 1995b , Inglot et al., (1997), Kallenberg and Ledwina, (1997), Kallenberg and Ledwina, (1999), Janic-Wróblewska and Ledwina, (2000), Kraus, (2007) and adopt the following finite-sample approximation for the null distribution of $N\hat{\Psi}_{\hat{K}}^{2}$:
@@
Such an approximation primarily accounts for the selection uncertainty of $\hat{K}$. In finite samples, under the null, the empirical probability of the event $\{\hat{K}=1\}$ may not be exactly $1$, as $\{\hat{K}=2\}$ can occur with a small but negligible probability. Consequently, (4.3) is derived from the approximation $\mathrm{P}(N\hat{\Psi}_{\hat{K}}^{2}\leq x)\approx\mathrm{P}(N\hat{\Psi}_{1}^{2}\leq x,\hat{K}=1)+\mathrm{P}(N\hat{\Psi}_{2}^{2}\leq x,\hat{K}=2)$. The reliability and accuracy of $H(x)$ are further illustrated through simulation studies in the next section and the appendix.
5 Simulation
In this section, we conduct Monte Carlo experiments to examine the finite-sample performance of the proposed tests developed in Sections 3 and 4. Following Neyman, (1937), Bera et al., (2013) and Song and Xiao, (2022), we adopt the orthonormal Legendre polynomials on the interval $[0,1]$ for $\{\pi_{k}\}_{k=0}^{\infty}$. Under this choice, the analytical properties of the constants $c_{1k}$ and $c_{2k}$ are provided in Proposition 1 of Duchesne et al., (2016), and numerical values for $1\leq k\leq 10$ are available in Table 1 of their Supplementary Material. From a computational perspective, the test statistic $N\hat{\Psi}_{K}^{2}$, introduced in Sections 3.1 and 3.2, can be computed using a simplified expression analogous to equation (14) in Duchesne et al., (2016), which leverages the closed form of $\bm{\Sigma}_{K}^{-1}$ [see equations (12) and (13) therein].
@@ 5 Simulation
Here $\{Y_{ij}^{(0,1)}\}_{i=1,j=1}^{jm,5}$ and $\{Y_{ij}^{(0,2)}\}_{i=1,j=1}^{jm,5}$ are generated under $H_{0}$, while $\{Y_{ij}^{(1,1)}\}_{i=1,j=1}^{jm,5}$ and $\{Y_{ij}^{(1,2)}\}_{i=1,j=1}^{jm,5}$ are generated under $H_{1}$. The sample size parameter $m$ varies from 10 to 150 in increments of 10. We set the significance level $\alpha=5\%$ and perform $500$ replications. We evaluate the test statistic $N\hat{\Psi}_{K}^{2}$ from Section 3.1, with $1\leq K\leq 5$, and the data-driven test statistic $N\hat{\Psi}_{\hat{K}}^{2}$ from Section 4, where $\hat{K}$ is determined by (4.1) with $D=5$.
@@
Results of Experiment I are shown in Tables 2, 3, 4 and Figure 1. Table 2 reports the empirical rejection rates under $H_{0}$. For fixed $K$, $N\hat{\Psi}_{K}^{2}$ maintains the nominal level well, even for small sample sizes such as $m=10$ (corresponding to a total sample size of $N=150$). For the data-driven test statistic $N\hat{\Psi}_{\hat{K}}^{2}$, the limiting $\chi^{2}_{1}$ distribution underestimates the variability under $H_{0}$, while the revised distribution $H(x)$ achieves better size control. Tables 3 and 4 present the empirical frequency of $\hat{K}$ under $H_{0}$ and $H_{1}$, respectively. Under $H_{0}$, the frequency of $\{\hat{K}=1\}$ approaches $1$ as $m$ increases, whereas under $H_{1}$, $\hat{K}$ is consistently greater than $1$, in line with theoretical expectations. Figure 1 displays the sample means of $\hat{K}$ across different values of $m$, with error bars representing one standard deviation around the mean. Under $H_{0}$, the sample mean of $\hat{K}$ stays close to $1$ with diminishing variability as $m$ grows, indicating consistent model selection. Under $H_{1}$, $\hat{K}$ increases steadily and approaches the maximum $D=5$, with reduced variability, demonstrating the adaptiveness of the data-driven procedure.

### v4

@@ 4 Data-driven choice of $K$
The methodology in Section 3 relies on the fact that the order of smooth alternatives (2.1), $K$, is fixed before testing. The choice of $K$ critically affects the performance of smooth tests (ledwina1994data; kallenberg1995consistency; kallenberg1995data; inglot1996asymptotic; inglot1997data; kallenberg1997dataA; kallenberg1997dataB; kallenberg1999data; janic2000data; ducharme2004smooth; ducharme2004goodness; kraus2007data; duchesne2016estimating). A large $K$ introduces redundant components into the alternatives that contribute little to the test statistic but inflate the degrees of freedom, thereby diluting power. Conversely, a small $K$ increases the risk that $\mathbb{E}[\pi_{k}(Z)]=0$ for all $1\leq k\leq K$, rendering the test powerless. Hence, an appropriate choice of $K$ is crucial. In this section, we focus on the data-driven selection of $K$ for the three tests developed in Section 3.
In Neyman’s smooth test literature, the data-driven selection of $K$ was first proposed by ledwina1994data for testing uniformity. The data-driven testing procedure consists of two steps. First, Schwarz’s selection rule (also known as the Bayesian Information Criterion, BIC) is applied to determine the dimension of the smooth model that best fits the data. Second, Neyman’s smooth test is performed within the selected model space, yielding a data-driven test statistic. The data-driven smooth test thus combines preliminary model selection with a more precise inferential procedure, serving as Neyman’s test in the “right” direction. Desirable theoretical properties and extensive numerical results of the selection rule and the induced test statistic were established in kallenberg1995consistency and kallenberg1995data, inglot1996asymptotic. Later, modifications of Schwarz’s rule addressed more complex problems, including testing composite hypotheses (kallenberg1997dataA; kallenberg1997dataB; inglot1997data), testing independence (kallenberg1999data), and two-sample testing (janic2000data), among others. These modified rules leverage smooth test statistics directly, avoiding the computation of the maximized log-likelihood, and are easier to implement.
@@
$\widehat{K}=\min\left\{\mathop{\arg\max}\limits_{1\leq k\leq D}\left(N\widehat{\Psi}_{k}^{2}-k\log N\right)\right\},$ (4.1)
where $D$ is a fixed positive integer, and $\widehat{\Psi}_{k}^{2}$ takes the form of (3.3) or (3.12) depending on the specific scenario. The resulting data-driven test statistic is $N\widehat{\Psi}_{\widehat{K}}^{2}$, with $\widehat{K}$ given by (4.1). The form of (4.1) is mainly inspired by ducharme2004smooth, ducharme2004goodness, kraus2007data and duchesne2016estimating where the upper bound of selection $D$ is fixed. An alternative is to let $D=D(N)\to\infty$ as $N\to\infty$ (see, for example, kallenberg1995consistency; kallenberg1995data; inglot1996asymptotic; inglot1997data; kallenberg1997dataA; kallenberg1997dataB; kallenberg1999data; janic2000data). Although a diverging upper bound can, in principle, improve the consistency of smooth tests, the required rate of divergence is slow, and simulation studies show that empirical power levels off rapidly as $D$ increases. Hence, our proposed selection rule (4.1) with fixed $D$ is both reasonable and easy to implement in practice.
To establish the theoretical properties of $\widehat{K}$ and $N\widehat{\Psi}_{\widehat{K}}^{2}$, we introduce a revised version of the alternative hypothesis $H_{1}$ (3.4) and Assumption 3.2. Consider
$H_{1}^{\prime}:\mathbb{E}[\pi_{1}(Z)]=\ldots=\mathbb{E}[\pi_{K_{0}-1}(Z)]=0,\mathbb{E}[\pi_{K_{0}}(Z)]\neq 0,\quad K_{0}\leq D,$ (4.2)
@@
    For model (3.1) in Section 3.1, as $N\to\infty$, under $H_{0}:e\sim\Phi(x)$, $\mathbb{P}(\widehat{K}=1)\to 1$ and $N\widehat{\Psi}_{\widehat{K}}^{2}\overset{d}{\to}\chi_{1}^{2}$; under $H_{1}^{\prime}$ in (4.2), $\mathbb{P}(\widehat{K}\geq K_{0})\to 1$ and $\mathbb{P}(N\widehat{\Psi}^{2}_{\widehat{K}}\leq x)\to 0$ for any $x\in\mathbb{R}$.
@@
    For model (3.9) in Section 3.2, as $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$, and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$, under $H_{0}:e\sim\Phi(x)$, $\mathbb{P}(\widehat{K}=1)\to 1$ and $N\widehat{\Psi}_{\widehat{K}}^{2}\overset{d}{\to}\chi_{1}^{2}$; under $H_{1}^{\prime}$ in (4.2), $\mathbb{P}(\widehat{K}\geq K_{0})\to 1$ and $\mathbb{P}(N\widehat{\Psi}^{2}_{\widehat{K}}\leq x)\to 0$ for any $x\in\mathbb{R}$.
@@
    For model (3.10) in Section 3.3, with additional conditions such that (i) there exist $0<\underline{\sigma}\leq\overline{\sigma}<\infty$ such that $\underline{\sigma}<\inf_{1\leq j\leq J}\sigma_{j}\leq\sup_{1\leq j\leq J}\sigma_{j}<\overline{\sigma}$; (ii) there exist $0<\underline{q}\leq\overline{q}<\infty$ such that $\underline{q}<\inf_{1\leq j\leq J}q_{j}\leq\sup_{1\leq j\leq J}q_{j}<\overline{q}$, as $\min\{N_{1},\ldots,N_{J}\}\to\infty$ and $J=o(N^{1/2})$, under $H_{0}:e\sim\Phi(x)$, $\mathbb{P}(\widehat{K}=1)\to 1$, $N\widehat{\Psi}_{\widehat{K}}^{2}\overset{d}{\to}\chi_{1}^{2}$ and $N\widetilde{\Psi}_{\widehat{K}}^{2}\overset{d}{\to}\chi_{1}^{2}$; under $H_{1}^{\prime}$ in (4.2), $\mathbb{P}(\widehat{K}\geq K_{0})\to 1$ and $\mathbb{P}(N\widetilde{\Psi}^{2}_{\widehat{K}}\leq x)\to 0$ for any $x\in\mathbb{R}$.
Theorem 10 presents the unified results of data-driven smooth tests for the three cases in Section 3. Under the null hypothesis, the probability of $\{\widehat{K}=1\}$ tends to $1$ asymptotically, implying that the first-order smooth model provides the best fit to $\{\widehat{Z}_{ij}\}_{i=1,j=1}^{N_{j},J}$ among the $D$ candidate models. From the perspective of hypothesis testing, this means that $\widehat{\Psi}_{1}^{2}$ is informative and sufficient for assessing uniformity or normality. Meanwhile, the data-driven selection procedure allows identification of the mechanism underlying nonuniformity (equivalently, nonnormality) and enhances the test’s power against a broader class of alternatives, $H_{1}^{\prime}$. These findings are consistent with previous results on data-driven smooth tests.
In practice, however, the $\chi_{1}^{2}$ limiting null distribution of data-driven smooth test statistics often performs poorly in finite samples (kallenberg1995data; inglot1997data; kallenberg1997dataA; kallenberg1997dataB; kallenberg1999data; janic2000data; ducharme2004smooth; ducharme2004goodness; kraus2007data; duchesne2016estimating). To address this issue in our problem, and provided the nested orthonormal system (i.e. $\{\pi_{k}\}_{k=1}^{K}\subseteq\{\pi_{k}\}_{k=1}^{K+1}$ for $1\leq K\leq D-1$), we follow the approach of kallenberg1995data, inglot1997data, kallenberg1997dataA, kallenberg1999data, janic2000data, kraus2007data and adopt the following finite-sample approximation for the null distribution of $N\widehat{\Psi}_{\widehat{K}}^{2}$:
@@
Such an approximation primarily accounts for the selection uncertainty of $\widehat{K}$. In finite samples, under the null, the empirical probability of the event $\{\widehat{K}=1\}$ may not be exactly $1$, as $\{\widehat{K}=2\}$ can occur with a small but non-negligible probability. Consequently, (4.3) is derived from the approximation $\mathbb{P}(N\widehat{\Psi}_{\widehat{K}}^{2}\leq x)\approx\mathbb{P}(N\widehat{\Psi}_{1}^{2}\leq x,\widehat{K}=1)+\mathbb{P}(N\widehat{\Psi}_{2}^{2}\leq x,\widehat{K}=2)$. The reliability and accuracy of $H(x)$ are further illustrated through simulation studies in the next section and in the Online Appendix.
5 Simulations
In this section, we conduct Monte Carlo experiments to examine the finite-sample performance of the smooth tests developed in Sections 3 and 4. Following neyman1937smooth, bera2013smooth, song2022smooth, and beutner2025two, we adopt the orthonormal Legendre polynomials on the interval $[0,1]$ for $\{\pi_{k}\}_{k=1}^{\infty}$. Under this choice, the analytical properties of the constants $c_{1k}$ and $c_{2k}$ are provided in Proposition 1 of duchesne2016estimating, and numerical values for $1\leq k\leq 10$ are available in Table 1 of their Supplementary Material. From a computational perspective, the test statistic $N\widehat{\Psi}_{K}^{2}$, introduced in Sections 3.1 and 3.2, can be computed using a simplified expression analogous to equation (14) in duchesne2016estimating, which leverages the closed form of $\bm{\Sigma}_{K}^{-1}$ (see equations (12) and (13) therein). For comparison, we include three classical normalty tests in the literature—Shapiro–Wilk (SW), Jarque–Bera (JB), and Kolmogorov–Smirnov (KS) tests—applied to the standardized residuals $\{\widehat{e}_{ij}\}_{i=1,j=1}^{N_{j},J}$ under the corresponding model, without accounting for the parameter estimation effects.
@@ 5 Simulations
Here $\{Y_{ij}^{(0,1)}\}_{i=1,j=1}^{jm,5}$ and $\{Y_{ij}^{(0,2)}\}_{i=1,j=1}^{jm,5}$ are generated under $H_{0}$, while $\{Y_{ij}^{(1,1)}\}_{i=1,j=1}^{jm,5}$ and $\{Y_{ij}^{(1,2)}\}_{i=1,j=1}^{jm,5}$ are generated under $H_{1}$. For the proposed smooth tests, we evaluate the test statistic $N\widehat{\Psi}_{K}^{2}$ from Section 3.1 with $1\leq K\leq 5$, as well as the data-driven statistic $N\widehat{\Psi}_{\widehat{K}}^{2}$ from Section 4, where $\widehat{K}$ is determined by (4.1) with $D=5$. The sample size parameter $m$ ranges from 10 to 150 in increments of 10, and the significance level is set at $\alpha=5\%$ with $500$ replications performed.
@@
Results of Experiment I are presented in Tables 2, 3, and Figure 1. Table 2 reports the empirical rejection rates under $H_{0}$ for the proposed smooth tests and the classical normality tests. The smooth tests comprise both fixed-$K$ versions ($K=1,\ldots,5$) and data-driven procedures, implemented using the limiting $\chi^{2}_{1}$ null distribution ($\widehat{K}\&\chi^{2}_{1}$) or the approximated null distribution $H(x)$ ($\widehat{K}\&H(x)$). For fixed $K$, under $H_{0}$, the statistic $N\widehat{\Psi}_{K}^{2}$ maintains the nominal level well, even for small sample sizes such as $m=10$ (corresponding to a total sample size of $N=150$). For the data-driven test statistic $N\widehat{\Psi}_{\widehat{K}}^{2}$, the limiting $\chi^{2}_{1}$ distribution fails to control the rejection rate under $H_{0}$, whereas the approximated distribution $H(x)$ achieves better size control, yielding rejection rates close to the nominal $5\%$ level. Under $H_{1}$, all versions of the proposed smooth tests exhibit high power across the considered settings. As for the classical methods, although the Shapiro–Wilk and Jarque–Bera tests are applied directly to the standardized residuals without accounting for the estimation effects, their empirical rejection rates under $H_{0}$ remain close to the nominal $5\%$ level. Under $H_{1}$, both tests achieve high power, indicating that they still provide reasonably reliable numerical performance despite the lack of formal adjustment for the estimation effect. In contrast, under $H_{0}$, the Kolmogorov–Smirnov test exhibits almost zero rejection across all settings, indicating the practical failure of the test. Table 3 presents the empirical frequency of $\widehat{K}$ under $H_{0}$ and $H_{1}$, respectively. Under $H_{0}$, the frequency of $\{\widehat{K}=1\}$ approaches $1$ as $m$ increases, whereas under $H_{1}$, $\widehat{K}$ is consistently greater than $1$, in line with theoretical expectations. Figure 1 displays the sample means of $\widehat{K}$ across different values of $m$, with error bars representing one sample standard deviation around the sample means. Under $H_{0}$, the sample mean of $\widehat{K}$ stays close to $1$ with diminishing variability as $m$ grows, indicating consistent model selection. Under $H_{1}$, $\widehat{K}$ increases steadily and approaches the maximum $D=5$, with reduced variability, demonstrating the adaptiveness of the data-driven procedure.